# Efficient optimization of neural network backflow for ab-initio quantum chemistry

### v1

@@ Abstract
The ground state of second-quantized quantum chemistry Hamiltonians is key to determining molecular properties. Neural quantum states (NQS) offer flexible and expressive wavefunction ansatze for this task but face two main challenges: highly peaked ground-state wavefunctions hinder efficient sampling, and local energy evaluations scale quartically with system size, incurring significant computational costs. In this work, we develop algorithmic improvements for optimizing these wave-functions which includes compact subspace construction, truncated local energy evaluations, improved stochastic sampling, and physics-informed modifications. We apply these improvements to the neural network backflow (NNBF) ansatz finding that they lead to improved accuracy and scalability. Using these techniques, we find NNBF surpasses traditional methods like CCSD and CCSD(T), outperform existing NQS approaches, and achieve competitive energies compared to state-of-the-art quantum chemistry methods such as HCI, ASCI, and FCIQMC. An ablation study highlights the contribution of each enhancement, showing significant gains in energy accuracy and computational efficiency. We also examine the dependence of NNBF expressiveness on the inverse participation ratio (IPR), observing that more delocalized states are generally harder to approximate.
@@ I Introduction
In this work, building upon our previous study [34], we introduce several algorithmic advancements which significantly improve the scalability and accuracy of general NQS optimization in QC. These enhancements include an efficient method for constructing a compact yet important subspace, a truncated local energy evaluation, an improved stochastic sampling method, and the incorporation of prior physical knowledge into the ansatz and training pipeline. These techniques build a general framework for improved optimization of second-quantized QC Hamiltonians with NQS. Using these techniques, we optimize the NNBF ansatz across various molecular systems—including the paradigmatic strongly correlated N2 molecule—demonstrating that NNBF not only continues to outperform traditional methods like CCSD and CCSD(T) and all other existing NQS approaches but also achieves competitive performance against state-of-the-art QC methods such as HCI [9, 10], ASCI [11, 12], and FCIQMC [38]. To quantify the contribution of each proposed improvement, we conduct an ablation study that cumulatively adds these techniques. The results show that our enhanced approach achieves orders-of-magnitude improvements in energy accuracy while substantially reducing the wall clock time per optimization step. Furthermore, we investigate how the expressiveness of the NNBF ansatz depends on the inverse participation ratio (IPR) of the target quantum state, providing additional insights into the factors influencing the representational capacity of NQS.
@@ II Methods
Figure 1: (a) The entire VMC workflow can be broadly divided into four components: (1) the total space $\mathcal{U}$, where amplitudes are computed at each iteration; (2) the sample $\mathcal{S}(\mathcal{U})$ drawn from $\mathcal{U}$; (3) the importance weights $w_{i}(\mathcal{S})$ assigned to the sampled configurations; and (4) the computation of the local energy $E_{l}(\mathbf{x}|\mathcal{A})$ given a subspace $\mathcal{A}$, where $\mathcal{A}$ denotes the set of configurations considered when iterating over all single and double excitations |$\ket{\mathbf{x}_{j}}\in\{D^{SD}_{i}\}$ Examples of $\mathcal{A}$ are the full Hilbert space $\mathcal{H}$ or the preselected subspace $\tilde{\mathcal{U}}$ (see part (c)) With these four components specified, both the energy and its gradient can be approximated and subsequently used to update the model parameters. (b) Workflow of fixed-size selected configuration (FSSC) [34]. The total space is the union of the core and connected spaces, i.e., $\mathcal{U}=\mathcal{V}\cup\mathcal{C}$. The sample is deterministically constructed by selecting the $k$ unique configurations with the largest amplitude magnitudes, and the importance weights are probabilities renormalized with respect to this sample. The local energy is computed exactly. (c) In this work, we preselect the connected space $\mathcal{C}$ via three steps represented by the three circles. First we identify all connections (via $H$) between core configurations |$\ket{\mathbf{x}_{i}}\in\mathcal{V}$ and connected configurations |$\ket{\mathbf{x}_{j}}\in\mathcal{C}$ and compute their strength $H_{ji}\psi_{\theta}(\mathbf{x}_{i})$ represented by weighted edges. Then (second circle) only the edge with largest weight is kept for each configuration |$\ket{\mathbf{x}_{j}}$. Finally (third circle) the configurations |$\ket{\mathbf{x}_{j}}$ with the largest $k^{\prime}$ values as well as all of $\mathcal{V}$ make up the pre-selected subspace $\tilde{\mathcal{U}}$. Amplitudes for this restricted subspace $\tilde{\mathcal{U}}$ are computed to define new probabilities $\tilde{p}_{i}$, which are used for sampling and streamlining local energy calculations via lookup: $\tilde{E}_{l}(\mathbf{x}_{i}|\tilde{\mathcal{U}})$. (d) The sample $\tilde{\mathcal{S}}$ is stochastically generated using the Gumbel top-$k$ trick. By perturbing the log-probabilities $\tilde{p}_{i}$ with Gumbel noise $g_{i}\overset{i.i.d.}{\sim}\text{Gumbel}(0,1)\quad\text{with pdf}\quad f(x)=e^{-(x+e^{-x})}$ the top-$k$ configurations with the largest perturbed log-probabilities are selected to form the sample $\tilde{\mathcal{S}}$. The associated weights $\tilde{w}_{i}(\kappa)$ provide unbiased estimates over $\tilde{\mathcal{U}}$, yielding improved approximations for the full Hilbert space.
@@ II.1 Overview
For a many-body system containing $N_{e}$ electrons and $N_{o}(>N_{e})$ spin-orbitals =$\mathcal{B}=\{\phi_{i}\}_{i=1}^{N_{o}}$, the many-electron wavefunction can be expressed in second quantization as |$\ket{\psi}=\sum_{i}\psi(\mathbf{x}_{i})\ket{\mathbf{x}_{i}}$ where |$\ket{\mathbf{x}_{i}}=\ket{x_{i}^{1\uparrow},\dots,x_{i}^{N_{o}\uparrow},x_{i}^{1\downarrow},\dots,x_{i}^{N_{o}\downarrow}}$ is the i-th computational basis vector, with $x_{i}^{j\uparrow/\downarrow}\in\{0,1\}$ denoting the occupation of the j-th spin-orbital. NQS offer a promising framework for efficiently representing many-electron wavefunctions via machine learning architectures: |$\ket{\psi_{\theta}}=\sum_{i}\psi_{\theta}(\mathbf{x}_{i})\ket{\mathbf{x}_{i}}$ where $\theta$ is model parameter. Among various NQS architectures, NNBF has demonstrated strong performance in fermionic systems [23, 24, 22, 34]. The NNBF wavefunction is defined as $\psi_{\theta}(\mathbf{x}_{i})=\sum_{m=1}^{D}\det[\Phi^{m}_{j=\{l|x_{i}^{l}=1\},k}(\mathbf{x}_{i};\theta)]$ where $\Phi^{m}_{jk}$ are ”configuration-dependent” single-particle orbitals (SPOs) output by NNBF’s internal network.
@@
. The gradient of energy is then given by
@@
In this work, we build on the latter approach. For each optimization step, we start from a core space $\mathcal{V}$ containing a fixed set of unique, important configurations. We then introduce a method to construct a larger important subspace $\tilde{\mathcal{U}}$. From this larger subspace, we generate a more efficient approximation of the local energy $\tilde{E}_{l}(\mathbf{x}|\tilde{\mathcal{U}})$. Additionally, we develop an improved strategy for sampling configurations $\tilde{\mathcal{S}}$ and assigning importance weights $\tilde{w}$, resulting in more accurate estimates of equation (1) and (3). We also generate a new core space for the next optimization step. These are the critical steps required to improve the accuracy and efficiency of the optimization.
II.2 Preselection
In this subsection, we outline our method for generating a compact yet dominant subspace $\tilde{\mathcal{U}}$, which preselects important configurations used throughout the algorithm.
We define the connected space $\mathcal{C}$ as the set of all other configurations connected to the core space $\mathcal{V}$ via non-zero Hamiltonian elements. With a fixed core space size, $\mathcal{C}$ grows quartically with system size. Importantly, identifying configurations in $\mathcal{C}$ does not require amplitude evaluations.
Each connected configuration |$\ket{\mathbf{x}_{j}}$ can be linked to multiple core configurations |$\ket{\mathbf{x}_{i}}$, with connection strengths quantified by $\psi_{\theta}(\mathbf{x}_{i})H_{ij}$ where “strength” can be realized either through strong Hamiltonian connections ($H_{ij}$ large) or proximity to dominant configurations ($\psi_{\theta}(\mathbf{x}_{i})$ large). The strongest of these connections is then defined as the importance score for element $x_{j}$:
@@ II.2 Preselection
$I(\mathbf{x}_{j})=\underset{\ket{\mathbf{x}_{i}}\in\mathcal{V}}{max}\absolutevalue{\psi_{\theta}(\mathbf{x}_{i})H_{ij}}.$ | . (4)
This metric closely parallels the heuristic used in Heat-bath Configuration Interaction (HCI) [9] for determinant selection: $f_{HCI}(\mathbf{x_{j}})=\underset{i\in\{D^{core}\}}{max}(\absolutevalue{H_{ji}c_{i}})$, which approximates the first-order perturbative correction |$\absolutevalue{\frac{\sum_{i}H_{ji}c_{i}}{E_{0}-H_{jj}}}$ |.
To construct the pre-selected space $\tilde{\mathcal{U}}$, we retain the entire core space and select a fraction, set to $1/(N_{o}-N_{e})$ in this work, of $\mathcal{C}$ with the largest importance score:
@@
$\tilde{\mathcal{U}}=\mathcal{V}\cup\underset{\ket{\mathbf{x}_{j}}\in\mathcal{C}}{\text{argtop}k^{\prime}}[\underset{\ket{\mathbf{x}_{i}}\in\mathcal{V}}{\text{max}}|H_{ji}\psi_{\theta}(\mathbf{x}_{i})|]$ (5)
where $k^{\prime}=\absolutevalue{\mathcal{C}}/(N_{o}-N_{e})$. We define the probability distribution:
$\tilde{p}_{i}=\frac{\psi^{2}_{\theta}(\mathbf{x}_{i})}{\sum_{\ket{\mathbf{x}_{j}}\in\tilde{\mathcal{U}}}\psi^{2}_{\theta}(\mathbf{x}_{j})}\quad\forall\ket{\mathbf{x}_{i}}\in\tilde{\mathcal{U}},$ (6)
Notice that only amplitudes in $\tilde{\mathcal{U}}$ need to be computed and these amplitudes are stored for the rest of the optimization step, preventing resource waste on insignificant configurations.
We can better understand why these preselected configurations are a reasonable choice by recognizing that the connected space $\mathcal{C}$ naturally emerges during the local energy evaluation with respect to the core space $\mathcal{V}$:
$E_{\theta}\approx\frac{1}{\sum_{\ket{\mathbf{x}_{k}}\in\mathcal{V}}\psi^{2}_{\theta}(\mathbf{x}_{k})}\sum_{\ket{\mathbf{x}_{i}}\in\mathcal{V},\ket{\mathbf{x}_{j}}}\psi_{\theta}(\mathbf{x}_{i})H_{ij}\psi_{\theta}(\mathbf{x}_{j}).$ (7)
This expression shows that each connected configuration |$\ket{\mathbf{x}_{j}}$ may appear multiple times through different connections, contributing to the energy via terms weighted by $\psi_{\theta}(\mathbf{x}_{i})H_{ij}$. This observation highlights that computing $\psi_{\theta}(\mathbf{x}_{j})$ is only meaningful if at least one of these connection weighs is non-negligible. The importance score in equation (4) captures this idea by selecting the strongest connection for each |$\ket{\mathbf{x}_{j}}$, ensuring that configurations with negligible contributions are excluded from the preselected space.
@@ II.3 Streamlined local energy calculations
In this subsection, we propose a strategy, building on Section II.2, to accelerate the calculation of the local energy—identified as one of the primary computational bottlenecks in NQS-based quantum chemistry. Computing $E_{l}(\mathbf{x}_{i})=\sum_{\ket{\mathbf{x}_{j}}}H_{ij}\frac{\psi(\mathbf{x}_{i})}{\psi(\mathbf{x}_{j})}$ requires evaluating ansatz amplitudes for all |$\ket{\mathbf{x}_{j}}$ connected to |$\ket{\mathbf{x}_{i}}$ through nonzero $H_{ij}$. For second-quantized molecular Hamiltonians, the number of such terms grows quartically with system size, making this step computationally expensive.
To mitigate this cost, we introduce a truncated approximation of the local energy by leveraging the preselected important subspace $\tilde{\mathcal{U}}$ from Section II.2:
$\tilde{E}_{l}(\mathbf{x}_{i})=E_{l}(\mathbf{x}_{i}|\tilde{\mathcal{U}})=\sum_{\ket{\mathbf{x}_{j}}\in\tilde{\mathcal{U}}}H_{ij}\frac{\psi(\mathbf{x}_{j})}{\psi(\mathbf{x}_{i})}.$ (8)
Because the amplitudes for configurations in $\tilde{\mathcal{U}}$ are already precomputed, the evaluation of this truncated local energy avoids repeated, costly wavefunction evaluations. Storing $\tilde{\mathcal{U}}$ in lexicographical order further enables efficient amplitude retrieval through $O(\log(\absolutevalue{\tilde{\mathcal{U}}}))$ lookups.
Our approach for computing the local energy shares conceptual similarities with other recent methods [30, 32, 35, 36], where the local energy sum is restricted to the sample set $\mathcal{S}$ generated by their sampling schemes. However, a key distinction is that $\tilde{\mathcal{U}}$ in our method is both larger than $\mathcal{S}$ and prefiltered for importance. This ensures that the truncated sum retains significant contributions to the local energy while providing a more accurate approximation with comparable or even reduced computational cost.

### v2

@@ Abstract
The ground state of second-quantized quantum chemistry Hamiltonians is key to determining molecular properties. Neural quantum states (NQS) offer flexible and expressive wavefunction ansatze for this task but face two main challenges: highly peaked ground-state wavefunctions hinder efficient sampling, and local energy evaluations scale quartically with system size, incurring significant computational costs. In this work, we overcome these challenges by introducing a suite of algorithmic enhancements, which includes efficient periodic compact subspace construction, truncated local energy evaluations, improved stochastic sampling, and physics-informed modifications.
Applying these techniques to the neural network backflow (NNBF) ansatz, we demonstrate significant gains in both accuracy and scalability. Our enhanced method surpasses traditional quantum chemistry methods like CCSD and CCSD(T), outperforms other NQS approaches, and achieves competitive energies with state-of-the-art ab initio techniques such as HCI, ASCI, FCIQMC, and DMRG. A series of ablation and comparative studies quantifies the contribution of each enhancement to the observed improvements in accuracy and efficiency. Furthermore, we investigate the representational capacity of the ansatz, finding that its performance correlates with the inverse participation ratio (IPR), with more delocalized states being more challenging to approximate.
@@ I Introduction
In this work, building upon our previous study [34], we introduce a suite of algorithmic advancements that significantly improve the accuracy and scalability of Neural Quantum State (NQS) methods for quantum chemistry. These enhancements, which form a general and robust optimization framework, consist of four main components: an efficient method for constructing and periodically updating a compact yet important subspace; a truncated local energy evaluation strategy reusing pre-computed information; an improved stochastic sampling method to provide more unbiased energy estimations; and the incorporation of prior physical knowledge into the ansatz and training pipeline.
We apply this framework to the NNBF ansatz and benchmark its performance on various challenging systems, including the paradigmatic strongly correlated N2 molecule. Our results show that the enhanced NNBF method not only continues to outperform traditional methods like CCSD and CCSD(T) and all other existing NQS approaches but also achieves competitive performance against state-of-the-art ab-initio methods such as HCI [9, 10], ASCI [11, 12], FCIQMC [38], and DMRG [39]. To quantify the contribution of each proposed improvement, we conduct an ablation study that cumulatively adds these techniques. The results show that our enhanced approach achieves orders-of-magnitude improvements in energy accuracy while substantially reducing the wall clock time per optimization step. To provide deeper insight, this is followed by two focused studies: one dissecting the components of our stochastic sampling method, and another comparing our local energy evaluation strategy against common alternatives. Finally, we investigate how the expressiveness of the NNBF ansatz depends on the inverse participation ratio (IPR) of the target quantum state, providing additional insights into the factors influencing the representational capacity of NQS.
@@ II Methods
Figure 1: (a) The Variational Monte Carlo (VMC) workflow. Each optimization iteration consists of four subroutines: (upper right) Identification (every $l$ iterations) of configuration in the target space $\mathcal{U}$ within the full Hilbert space $\mathcal{H}$; a sample set $\mathcal{S}$ will be drawn from $\mathcal{U}$ at each optimization step and corresponding importance weights $w_{i}$ are assigned; (lower right) The local energies $E_{l}(\mathbf{x}_{i}|\mathcal{A})$ are computed for each sample $\mathbf{x}_{i}\in\mathcal{S}$ given a subspace $\mathcal{A}$, where $\mathcal{A}$ restricts the set of configurations considered when iterating over all connected configurations expanded from |$\ket{\mathbf{x}_{i}}$. Examples of $\mathcal{A}$ are the full Hilbert space $\mathcal{H}$ (when computed exactly) or the target space $\mathcal{U}$; (lower left) The total energy and its gradient are estimated via an importance-sampled average using the samples $\mathcal{S}$, weights $w_{i}$, and local energies $E_{l}$. (upper left) An optimizer, such as gradient descent or Stochastic Reconfiguration, uses these estimates to update the model parameters. (b) The Intermittent Target Selection (ITS) update, performed every $l$ steps. First, the core space $\mathcal{V}$ is updated by selecting the |$\absolutevalue{\mathcal{V}}$ | highest-amplitude unique configurations from the union of the prior target space $\mathcal{U}$ and a set of |$\absolutevalue{\mathcal{V}}$ | MCMC walkers maintained concurrently to provide access to the current wave-function. Subsequently, a new target subspace $\mathcal{U}$ is constructed by selecting the |$\absolutevalue{\mathcal{C}}/l$ configurations with largest amplitude moduli from the updated core space and its newly generated connected space. This subspace $\mathcal{U}$ then remains fixed until hitting the next ITS step, with all amplitude calculations being restricted to this compact yet important subspace $\mathcal{U}$. (c) Workflow of the Fixed-Size Selected Configuration (FSSC) method [34] for comparison. The target space is the union of the core and connected spaces, i.e., $\mathcal{U}=\mathcal{V}\cup\mathcal{C}$. The sample is deterministically constructed by selecting the $k$ unique configurations with the largest amplitude magnitudes, and the importance weights are probabilities renormalized with respect to this sample. The local energy is computed exactly. (d) Workflow for the stochastic sampling via the Gumbel top-$k$ trick used in the current work . The log-probabilities $\log{p_{i}}$ ) of configurations in the target space $\mathcal{U}$ are perturbed with Gumbel noise $g_{i}\overset{i.i.d.}{\sim}\text{Gumbel}(0,1)\quad\text{with pdf}\quad f(x)=e^{-(x+e^{-x})}$. The sample $\mathcal{S}$ is then formed by selecting the $k$ configurations with the largest perturbed log-probabilities. The corresponding importance weights are calculated to provide unbiased estimates over the subspace $\mathcal{U}$, using a threshold $\kappa$ defined as the value of the $(k+1)$-th largest perturbed log-probability.
@@ II.1 Overview
For a many-body system containing $N_{e}$ electrons and $N_{o}$ single-particle orbitals (SPOs) =$\mathcal{B}=\{\phi_{i}\}_{i=1}^{N_{o}}$, the many-electron wavefunction can be expressed in second quantization as |$\ket{\psi}=\sum_{i}\psi(\mathbf{x}_{i})\ket{\mathbf{x}_{i}}$ where |$\ket{\mathbf{x}_{i}}=\ket{x_{i}^{1\uparrow},\dots,x_{i}^{N_{o}\uparrow},x_{i}^{1\downarrow},\dots,x_{i}^{N_{o}\downarrow}}$ is the i-th computational basis vector, with $x_{i}^{j}\in\{0,1\}$ denoting the occupation of the j-th spin-orbital. NQS offer a promising framework for efficiently representing many-electron wavefunctions via machine learning architectures: |$\ket{\psi_{\theta}}=\sum_{i}\psi_{\theta}(\mathbf{x}_{i})\ket{\mathbf{x}_{i}}$ where $\theta$ are model parameters. Among various NQS architectures, NNBF has demonstrated strong performance in fermionic systems [23, 24, 22, 34]. The NNBF wavefunction is defined as $\psi_{\theta}(\mathbf{x}_{i})=\sum_{m=1}^{D}\det[\Phi^{m}_{j=\{l|x_{i}^{l}=1\},k}(\mathbf{x}_{i};\theta)]$ where $\Phi^{m}_{jk}$ are “configuration-dependent” spin-orbitals output by NNBF’s internal network.
@@
where indices $i,j,k,l$ iterate over the $N_{o}$ SPOs and $\sigma,\sigma^{\prime}$ denote spin. The gradient of energy is then given by
@@
In this work, we build on the latter, SCI-inspired approach. There is a core space $\mathcal{V}$, containing a set of unique and dominant configurations, and a target space $\mathcal{U}$, containing a larger compact yet relevant set of important determinants, both of which are updated every $l$ optimization steps. This periodic update takes as input the current $\mathcal{U}$ as well as a set of walkers generated stochastically from MCMC. From $\mathcal{U}$, we generate a more efficient and effective approximation of the local energy $E_{l}(\mathbf{x}|\mathcal{U})$. Additionally, we develop an improved strategy for sampling configurations $\mathcal{S}$ and assigning importance weights $w$, resulting in more accurate estimates of equation (1) and (3). These are the critical steps required to improve the accuracy and efficiency of the optimization.
II.2 Intermittent Target Selection
In this subsection, we introduce Intermittent Target Selection (ITS), a method to construct and periodically update a compact yet highly relevant subspace, denoted as $\mathcal{U}$, every $l$ optimization steps. This subspace and its amplitudes, which are calculated at each step, are then used in subsequent optimization subroutines including sample drawing and local energy computations to improve computational efficiency.
The output from ITS is the new target space $\mathcal{U}$ with input as the $\mathcal{U}$ that was selected $l$ steps ago and has been fixed in the last $l$ steps. At this point in the algorithm, we have $\psi_{\theta}(\mathbf{x}_{i})$ for all |$\ket{\mathbf{x}_{i}}\in\mathcal{U}$ computed from the last optimization step. The first step of ITS is to re-generate a new core space $\mathcal{V}$ by choosing |$\absolutevalue{\mathcal{V}}$ | unique configurations with the highest amplitude moduli from both $\mathcal{U}$ as well as the |$\absolutevalue{\mathcal{V}}$ | samples generated from a parallel MCMC track, which gives access to stochastic samples from the current wave-function - i.e.
@@ II.2 Intermittent Target Selection
$\mathcal{V}\leftarrow\underset{\ket{\mathbf{x}_{i}}\in\mathcal{MC}\cup\mathcal{U}}{\text{argtop}(\absolutevalue{\mathcal{V}})}\left[|\psi_{\theta}(\mathbf{x}_{i})|\right]$ (4)
[…]
@@
$\mathcal{U}\leftarrow\underset{\ket{\mathbf{x}_{i}}\in\mathcal{V}\cup\mathcal{C}}{\text{argtop}\left(\frac{\absolutevalue{\mathcal{C}}}{l}\right)}\left[|\psi_{\theta}(\mathbf{x}_{i})|\right]$ (5)
[…]
This strategy effectively amortizes the high cost of exploring the full connected space. The average number of amplitude evaluations is reduced by a factor of approximately $l$ compared to a naive approach that evaluates the entire space $\mathcal{C}$ at every step. The interval $l$ is chosen to align this amortized cost with the per-step cost of amplitude computations within $\mathcal{U}$, which requires |$\absolutevalue{\mathcal{C}}/l$ NNBF evaluations; we typically use $l\sim(N_{o}-N_{e})$. To manage the computational cost of the MCMC evolution, the number of MCMC walkers, $N_{w}$, is typically set to equal the core space size ($N_{w}=\absolutevalue{\mathcal{V}}$ |), and each walker performs $N_{e}$ proposed hopping moves per optimization step.
The effectiveness of ITS hinges on the assumption that the wavefunction parameters, $\theta$, evolve slowly enough that the changes over $l$ steps are minor. This ensures that the subspace $\mathcal{U}$ selected at the beginning of a cycle remains a good approximation of the most dominant configurations for the entire interval. Such slow parameter evolution is typically achieved by using a small learning rate and is most prominent as the optimization converges and the gradients naturally diminish.
@@ II.3 Streamlined Local Energy Calculations
Building upon the Intermittent Target Selection (ITS) strategy detailed in Section II.2, this subsection introduces an approach to accelerate the calculation of the local energy—identified as one of the primary computational bottlenecks in NQS-based quantum chemistry. Computing $E_{l}(\mathbf{x}_{i})=\sum_{\ket{\mathbf{x}_{j}}}H_{ij}\frac{\psi(\mathbf{x}_{i})}{\psi(\mathbf{x}_{j})}$ requires evaluating ansatz amplitudes for all |$\ket{\mathbf{x}_{j}}$ connected to |$\ket{\mathbf{x}_{i}}$ through nonzero $H_{ij}$. For second-quantized molecular Hamiltonians, the number of such terms grows quartically with system size, making this step computationally expensive.
To mitigate this cost, we introduce a truncated approximation of the local energy by leveraging the selectively important subspace $\mathcal{U}$ from Section II.2:
$E_{l}(\mathbf{x}_{i})=E_{l}(\mathbf{x}_{i}|\mathcal{U})=\sum_{\ket{\mathbf{x}_{j}}\in\mathcal{U}}H_{ij}\frac{\psi(\mathbf{x}_{j})}{\psi(\mathbf{x}_{i})}.$ (6)
Because the amplitudes for configurations in $\mathcal{U}$ are already computed, the evaluation of this truncated local energy avoids repeated, costly wavefunction evaluations. Storing $\mathcal{U}$ in lexicographical order further enables efficient amplitude retrieval through $O(\log(\absolutevalue{\mathcal{U}}))$ lookups.
Our approach for computing the local energy shares conceptual similarities with other recent methods [32, 35, 36], where the local energy sum is restricted to the sample set $\mathcal{S}$ generated by their sampling schemes. However, a key distinction is that $\mathcal{U}$ in our method is both substantially larger than $\mathcal{S}$ and constructed for importance. This strategy ensures that the local energy calculation retains more significant contributions, leading to a more accurate approximation and improved training performance at a comparable computational cost. A direct comparison of these strategies is presented in Section III.2.3.